Rama pays RM18,300 on 15 October 2018 for a loan of RM18,000 made on a certain date.The simple interest rate is 3% per annum. Determine the term of the loan (in days) and the date of the loan.


1
Expert's answer
2021-05-07T09:05:52-0400

A = P(1 + r/n)

In the formula

  • A = Accrued amount
  • P = Principal amount
  • r = Annual nominal interest rate as a decimal
  • R = Annual nominal interest rate as a percent
  • r = R/100
  • n = number of compounding periods per unit of time
  • t = time in decimal years
  • I = Interest amount

"18300=18000(1+0.03t)"

"t=0.55 days"

The date of the loan was 16th october because it was half a day loan.